OpenStudy (anonymous):

integral of ( x^2-2x=1)/x dx

OpenStudy (anonymous):

No boundaries? If not - just find the antiderivative and add a constant C.

OpenStudy (anonymous):

x^(3)/3-x^(2)-x+C

OpenStudy (anonymous):

Can you show me how you got that?

OpenStudy (earthcitizen):

\[\int\limits (x^2-2 x-1)/x dx = 1/2 ((x-4) x-2 \ln(x))+c\]

OpenStudy (anonymous):

Oh, missed the /x, nvm my answer, hehe

OpenStudy (anonymous):

Its x^2-2x+1

OpenStudy (anonymous):

You wrote: "x^2-2x=1"

OpenStudy (anonymous):

Yes i know i meant to but a +

OpenStudy (turingtest):

\[\int x^ndx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (earthcitizen):

\[\int\limits (x^2-2 x+1)/x dx = 1/2 (x-4) x+\ln(x)+c\]